# Slow is Fast! Dissecting Ethereum’s Slow Liquidity Drain Scams

### v2

@@ 7.1 Data Collection
DEX orders. Exchange activities and profit patterns within PoolData are essential for computing the profit-related metrics and profit-generating metrics (§6.1). We collected all types of exchange activities (via smart contract events) of six DEXs. Those activities include liquidity deposit, withdraw, and trade (buy/sell) between two tokens in each pool. Considering the data collection process operates independently of PoolData extraction, we specifically filtered the orders to include only those relevant to the liquidity pools in PoolData.
We collected over 917M DEX activities across 319,166 liquidity pools (Table 1). It accounts for 82.35% of the total Ethereum trading volume historically [55]. The resulting dataset, referred to as DEXData, contains rows where each entry represents an activity by either the pool owner or an investor within a specific liquidity pool. The dataset includes activities recorded up to July 13, 2024, where all pools in PoolData have at least one month of DEX activity history.
Smart contract security features.  Token security-related features (such as buy/sell tax, trading lock, etc.) from the paired token’s smart contract are critical for applying the heuristic’s Honeypot Validator. We utilized the security API provided by GoPlus [43] and De.Fi [39] to collect the token’s security characteristics indicative of honeypot behavior, including buy/sell tax, sell restriction, and more.
We extracted security features from 302,675 distinct tokens (base and paired) across 319,166 Ethereum liquidity pools, noting that some tokens are associated with multiple pools. These features were organized into a dataset, referred to as SecurityData, where each row represents a distinct token along with its complete set of security-related features.
7.2 Applying Heuristic for SLID Detection
We described how we apply the heuristic to detect Ethereum’s SLID scams. We began with data preprocessing and extracted the required features for all the liquidity pools. We then conducted an in-depth analysis of the flagged SLID pools across multiple dimensions, including vitality, profit, and trend.
Data preprocessing.  We preprocessed the three datasets (PoolData, DEXData, and SecurityData) to enhance the feature set and prepare for the heuristic detection phase.
For each liquidity pool instance in PoolData, we extracted all related DEX activities from DEXData to compile its complete activity history. Profit-related and owner-activity metrics were then computed (equations in Appendix C) and added as new features to the PoolData dataset. These enriched metrics serve as key inputs for the flagging process. Then, smart contract security features from SecurityData were mapped to their corresponding liquidity pools in PoolData.
Our preprocessing resulted in an updated dataset, called ExtendedPoolData, containing a comprehensive range of features (Table 4) required for heuristic detection.
@@ 7.2 Applying Heuristic for SLID Detection
    (Quick access to our dataset) Out of 319,166 selected liquidity pools, we flagged 3,117 instances of SLID scams. For evaluation, we randomly selected 100 SLID pools for manual observation, which revealed no sign of false positives. To validate the novelty of this scam tactic, we tested the randomly selected pools using existing honeypot and rug pull heuristics, none of which detected the scam.
(i) Vitality analysis of SLID pools. Figure 8 depicts the SLID liquidity pool age and the remaining active pool numbers in each age group. Of the 3,117 SLID liquidity pools, 70.8% (2,207/3,117) remained active after the first month, and 53.3% (1,660/3,117) were still operational three months after their deployment. Five pools reached the longest recorded age of four years. Among the 196 pools that were active for over three years, 28 are still alive today. The performance is exceptional compared to other scams (rug pull and honeypot). Over their lifetime, SLID pools attracted a total of 1.49M activities on DEXs, with a cumulative volume of \$4.49B.
We concluded that the SLID scams exhibit the characteristics of a long pool lifespan, with remarkably extended periods of activity and significant user engagement. This result can be further supported by their total trading volume.
Figure 6: Lifetime analysis of captured SLID liquidity pools.
@@
Figure 8: Daily volume (in USD) and activity count in SLID scam liquidity pools.
(ii) Profit analysis of SLID pools. We analysed the financial earnings of scammers. Across all flagged SLID liquidity pools, the scammers have lured 167,632 victims, generating \$79.1M in profit through a gradual profit-taking strategy. Additionally, unrealized profit assets worth \$24.1M remain under the control of scammers in active pools. On average, scammers achieved a profit of 6.08 times their initial investment. The top liquidity pools generating the highest profit are described in Table 2. The 3,117 SLID scams were initiated by 2,951 distinct addresses, with 90 addresses involved in more than one scam and one address deploying a peak of 23 SLID scams (details in Table 3).
We further investigated the scammers’ gradual strategy by analyzing their profit-taking patterns over time from the first day of pool deployment. Figure 8 illustrates the total number of profit-taking actions and the corresponding realized profit by days from each pool’s respective deployment date. We observed that most profit-taking actions occur in the first 100 days of the SLID liquidity pools, with 35.71% (31,912/89,364) of sell or withdraw actions executed during this period. While the number of profit-taking actions decreases over time, the total realized profit remains relatively stable throughout the evaluation period, ranging between \$1K and \$100K, with a few outliers.
@@
(iii) Trend analysis of SLID pools. We analysed the trend of collected pools. Figure 9 illustrates a steady increase in the number of SLID pools created annually, with the exception of 2024, where the lower number is due to the limited capture period (6 months). Meanwhile, Figure 8 illustrates the total activities and volume associated with SLID scams. We describe the development of SLID scams as follows.
@@
    Early development phase (2018 - 2020): During this period, a small number of SLID pools were created, exhibiting low volume and minimal activity. This could be attributed to the limited presence of DEXs and the lack of user familiarity with decentralized activities during this time [56]. However, there was a gradual increase in the development of SLID LPs, with recorded increments in user engagement.
@@
    Exploding phase (Mid-2020 - 2023): Following the rapid growth of DeFi, the market experienced a significant surge in SLID scams. The number of SLID liquidity pools created in 2020 was over 10 times greater than in 2019, accompanied by an explosion in trading volume and activity within these pools, with volumes ranging between \$10K and \$100M. This phase also saw record-breaking performances for SLID scams, with an all-time peak total volume of \$85.7M on March 8, 2021, and peak DEX activity counts of 17,586 on December 20, 2022, across all SLID LPs.
@@
    Stabilized phase (2023 - present): The statistics for SLID liquidity pools maintained high levels of stability, with daily activity counts ranging from 100 to 1000 and daily volumes ranging from \$10K to \$10M. During this period, the average daily volume was \$608,159, and there were 266.67 DEX activities per day across SLID pools. Although not exhibiting the same rapid growth as the previous one, the scam tactic remains attractive to scammers, as evidenced by the creation of 1,107 new pools. This continued appeal is likely due to the lack of prior recognition.
We conclude that there is a clear upward trend in the initialization of SLID scams over the years. The SLID strategy has been increasingly adopted by malicious actors.
7.3 ML Approach for Early SLID Warning
We take a step further. Our heuristic can only flag SLID pools only after profit-taking actions that trigger the heuristic’s validators. However, investors already lost their money.
We thus propose a proactive ML-based approach to raise early warnings and prevent users from engaging at the pool’s early stages. We utilized machine learning methods to enable fast detection. In particular, we extracted features from the liquidity pool in progressively shorter time windows to identify a reasonable timeframe that offers high prediction accuracy.
2018201920202021202220232024005005005005001,00010001{,}0001 , 0006464646452525252585585585585446446446446863863863863979979979979128128128128Year (p.s. 2024 up to June)SLID LPs
@@ 7.3 ML Approach for Early SLID Warning
Data preprocessing and feature extraction.  During our heuristic detection phase, we labeled all the collected liquidity pools. We still utilized these labels as ground truth for training the ML algorithms. It contains 3,117 flagged SLID pools and 316,049 non-SLID pools in the ExtendedPoolData dataset.
For early detection, we constrained the liquidity data (i.e., DEX activities) used for feature extraction to the first $d$ days after the deployment of pools in ExtendedPoolData.
During the preprocessing phase, we extracted features for each ground truth liquidity pool within the first $d$ days of its creation and appended these features to the corresponding liquidity pool records in ExtendedPoolData.
We then utilized the resulting dataset, called $d^{th}$-day, to train and evaluate ML algorithms. We varied the $d$ value, training the algorithms separately for each duration. We determine the shortest time frame by systematically reducing $d$, which balances high accuracy in detecting SLID scams with early warning capabilities. For pools with a lifetime shorter than $d$, features were extracted based on all available DEX activities from deployment to their inactive state.
To build an effective classifier within the specified time constraints, we extracted 57 features from DEXData, capturing diverse patterns of DEX activities. These features were categorized into four groups: Owner Activity Features (OAF), User Activity Features (UAF), Liquidity Pool Features (LPF), and Profit Features (PF) (details in Appendix G).
@@
Based on these considerations, we focused on three ensemble machine learning models: Random Forest [57], XGBoost [58], and Logistic Regression [59]. We exclude several alternative approaches, such as graph-based or network-based strategies, due to their high complexity and computational demands. Moreover, they are unsuitable for the scenarios we address [60] (more rationales in Appendix E).
@@
We selected the initial value $d$ as 267 days because, within this time frame, all four detectors remain to demonstrate optimal performance in SLID detection.
Detection results (Figure LABEL:fig:detection-metrics). As the time range was reduced, the performance of recall of the heuristic began to decline due to the limited number of order activities, which caused the heuristic validators to fail to trigger some detections. Given that the true negative count in the ground truth (316,049) is substantial compared to the true positive (3,117), both accuracy and precision of the heuristic remained unchanged when $d$ decreased. In contrast, the three ML algorithms maintained stable and good performances with decreasing time ranges. However, Random Forest performs consistently better than the other two ML algorithms in most $d$.
Random Forest was the last to maintain an accuracy greater than 93% down to $d$ = 56, whereas Logistic Regression and XGBoost saw their accuracy drop to 93% when using data within a time range of 58 days from the liquidity pool’s deployment date. Starting from $d$ = 56, the accuracy of all ML algorithms fell to 93%, marking the threshold at which all algorithms lost their high performance observed previously.
@@
8 Discussions
8.1 SLID Scam Mitigation
@@
    Assess on-chain data. Prioritize transparency and legitimacy by thoroughly evaluating on-chain activities. Regularly monitor the token to detect any signs of the owner gradually draining the liquidity pool.
8.2 SLID Limitations and Our Plan
Exceptional ambiguous cases. Initial LP token burning is an industry standard for ensuring pool transparency and preventing the owner from controlling 100% of the pool’s funds. However, this study does not cover cases where liquidity pools are created without burning LP tokens, and the owner takes profits only a few times, appearing legitimate. Such cases may initially resemble SLID patterns but later gain legitimacy as the token becomes popular (e.g., with meme coins or community-driven tokens).
These pools operate on the grey zone of the boundary between fraudulent activity and legitimate trading strategies while containing all fraudulent traits (i.e., the owner controls the pools with all investor’s money, can take them all anytime, and can do so multiple times through small profit-taking activities) but gain community approval. Those ambiguous cases make them challenging for us to classify as fraudulent confidently. Although the evaluation didn’t show the result of any involved legitimate liquidity pool being falsely labeled as SLID, we acknowledge this limitation that the heuristics (and ML models) might have limited ability in those cases and might result in false positives. Our future work aims to refine and update current heuristics to account for these ambiguous cases, potentially involving expanding features and community sentiment in heuristic labeling.
Advanced manipulations. Although our designed heuristic includes evaluations of the initial ground truth and legitimate tokens, we cannot yet address complex manipulations. We provide two example scenarios: (i) the liquidity pool conducts initial token burning for transparency and still attempts to commit with SLID actions; (ii) the pool owner does not conduct the action themselves but transfers LP tokens/inflated paired token amounts to another/others account(s) to execute a SLID strategy. We acknowledge that the true scope SLID may extend beyond what we detect, and we treat the present results as an upper bound and recognize the need for further refinements in the future.
8.3 Real-world Integration
SLID monitoring and early warning services. Our solution can be integrated into DEXs or crypto-related services as a LP monitoring service, a cross-checking mechanism, or a new security standard/requirement. A key requirement is the extraction of features, which can be directly derived from the token’s transaction history to evaluate LPs deployed on public blockchains, in particular, Ethereum. This service provides real-time warnings to investors about any fraudulent activities in existing or newly deployed tokens.
However, there are challenges. The dynamically required features for the ML models (e.g., volume and owner activity counts, which continuously change) demand significant computational resources to maintain and update. If a constant monitoring service is to be provided based on our solution, the model would require substantial resources to track and process these evolving features from the targeted pool. This limitation creates a barrier, making it nearly impossible to build an SLID detection service that can constantly update features and monitor all existing liquidity pools.
To date, the service can focus on a targeted set of suspect liquidity pools, enabling more effective resource allocation while still maintaining early detection.
Wallet credibility evaluation. Most payment provider services accept cryptocurrency transactions only from legitimate and well-established tokens. For instance, PayPal accepts cryptocurrency payments from decentralized wallets using Bitcoin and a few other valid tokens [63]). Newly created tokens are rarely considered by legitimate payment providers.
Integrating our solution with payment services enhances risk awareness in crypto-transactions, even if the SLID token is not yet listed. By identifying wallets linked to SLID creation as fraudulent and analyzing their transaction histories, our solution detects SLID activity and assesses fraud risks. This enables providers to flag fraudulent wallets, even when handling legitimate tokens like Bitcoin.
The limitation is the uncertainty of the required time to process transaction history to extract features. The computation time varies depending on the length of a pool’s historical transactions. One potential workaround is to provide offline detection. However, the continuous updating of pool transactions and their features presents a challenge.
9 Conclusion

### v3

@@ VI-B Large-scale SLID Detection
    (Quick access to our dataset). Out of our 319,166 selected LPs, we flagged 3,117 instances of SLID scams (accessed in [66]). We further conducted various steps to evaluate the reliability of the dataset (deferred to Appendix A-E):
(i) Vitality analysis of SLID pools. Figure 11 depicts the SLID LP age and the remaining active LP numbers in each age group. Of the 3,117 SLID LPs, 70.8% (2,207/3,117) remained active after the first month, and 53.3% (1,660/3,117) were still operational three months after their deployment. Five pools reached the longest recorded age of four years. Among the 196 pools that were active for over three years, 28 are still alive today. The performance is exceptional compared to other scams (rug pull and honeypot). Over their lifetime, SLID pools attracted a total of 1.49M activities on DEXs, with a cumulative volume of \$4.49B.
We concluded that SLID scams exhibit the characteristics of a long LP lifespan, with extended periods of activity and deep user engagement (further supported by trading volume).
Figure 9: Lifetime analysis of captured SLID LPs.
@@
Figure 11: Daily volume (in USD) and activity count in SLID scam LPs.
(ii) Profit analysis of SLID pools. We analyzed the financial earnings of scammers. Across all flagged SLID LPs, the scammers have lured 167,632 victims, generating \$79.1M in profit through a gradual profit-taking strategy. Additionally, unrealized profit assets worth \$24.1M remain under the control of scammers in active pools. On average, scammers achieved a profit of 6.08 times their initial investment. The top LPs generating the highest profit are described in Table II. The 3,117 SLID scams were initiated by 2,951 distinct addresses, with 90 addresses involved in more than one scam and one address deploying a peak of 23 SLID scams (Table III).
We further investigated the scammers’ gradual strategy by analyzing their profit-taking patterns from the first day of LP deployment. Figure 11 illustrates the total number of profit-taking actions and the corresponding realized profit by days from each pool’s respective deployment date. We observed that most profit-taking actions occur in the first 100 days of the SLID LPs, with 35.71% (31,912/89,364) of sell or withdraw actions executed during this period. While the number of profit-taking actions decreases, the total realized profit remains relatively stable throughout the evaluation period, ranging between \$1K and \$100K, with a few outliers.
@@
(iii) Trend analysis of SLID pools. We analysed the trend of collected pools. Figure 12 illustrates a steady increase in the number of SLID pools created annually, with the exception of 2024, where the lower number is due to the limited capture period (6 months). Figure 11 illustrates the total activities and volume associated with SLID scams. We describe the development of SLID scams below.
@@
    Early development phase (2018-2020): During this period, a small number of SLID pools were created, exhibiting low volume and minimal activity. This could be due to the limited presence of DEXs and the lack of user familiarity with decentralized activities [67]. However, there was a gradual increase in the development of SLID LPs, with recorded increments in user engagement.
@@
    Exploding phase (Mid-2020-2023): Following the rapid growth of DeFi, the market experienced a significant surge in SLID scams. The number of SLID LPs created in 2020 was over 10 times greater than in 2019, accompanied by an explosion in trading volume and activity within these pools, with volumes ranging between \$10K and \$100M. This phase also saw record-breaking performances for SLID scams, with an all-time peak total volume of \$85.7M on March 8, 2021, and peak DEX activity counts of 17,586 on December 20, 2022, across all SLID LPs.
@@
    Stabilized phase (2023-present): The statistics for SLID LPs maintained high levels of stability, with daily activity counts ranging from 100 to 1000 and daily volumes ranging from \$10K to \$10M. During this period, the average daily volume was \$608,159, and there were 266.67 DEX activities per day across SLID pools. Although not exhibiting the same rapid growth as the previous one, the scam tactic remains attractive to scammers, as evidenced by the creation of 1,107 new pools. This continued appeal is likely due to the lack of prior recognition.
[…]
VI-C ML Approach for Early SLID Warning
We take a step further. Our heuristic flags SLID pools only after profit-taking actions that trigger the heuristic’s validators. However, investors have already lost their money.
We thus propose a proactive, ML-based approach to raise early warnings and prevent users from engaging during the pool’s early stages. In addition, the ML model is also effective in identifying borderline SLID cases that may be missed by rule-based heuristics (detailed explanation in Appendix A-F1). In particular, we extracted features from the LP in progressively shorter time windows to identify a reasonable timeframe that offers high prediction accuracy.
201820192020202120222023202405005001,0001{,}00064645252585585446446863863979979128128Year (p.s. 2024 up to June)SLID LPs
@@ VI-C ML Approach for Early SLID Warning
Data preprocessing and feature extraction.  We labelled all LPs and utilised the labelled data for training ML algorithms. The dataset contains 3,117 flagged SLID pools and 316,049 non-SLID pools in the ExtendedPoolData dataset.
For early detection, we constrained the liquidity data (i.e., DEX activities) used for feature extraction to the first $d$ days after the deployment of pools in ExtendedPoolData. During the preprocessing phase, we extracted features for each LP within the first $d$ days of its creation and appended these features to the corresponding LP records in ExtendedPoolData.
We then utilized the resulting dataset, $d^{th}$-day, to train and evaluate ML algorithms. We varied the $d$ value, training the algorithms separately for each duration. We determine the shortest time frame by systematically reducing $d$, which balances high accuracy in detecting SLID scams with early warning capabilities. For pools with a lifetime shorter than $d$, features were extracted based on all available DEX activities from deployment to their inactive state.
To build an effective classifier within the specified timeframe, we extracted 57 features from DEXData, capturing patterns of DEX activities (Appendix A-H). These features were categorized as: Owner Activity Features (OAF), User Activity Features (UAF), LP Features (LPF), and Profit Features (PF).
26715010060595857560.850.850.90.90.950.9511
(a) Accuracy
26715010060595857560.850.850.90.90.950.9511
(b) Precision
26715010060595857560.850.850.90.90.950.9511
(c) Recall
26715010060595857560.850.850.90.90.950.9511
(d) F1 score
Figure 13: Detection evaluation metrics over varying days. Notations: X axis for days, Y axis for values; Heuristic in blue (first bar), Random Forest in red (second), Logistic Regression in brown (third), and XGBoost in black (fourth).
@@
Based on these considerations, we focused on three ML models: Random Forest [68], XGBoost [69], and Logistic Regression [70]. We exclude several alternative approaches, such as graph-based or network-based strategies, due to their high complexity and computational demands [71, 72]. Moreover, they are unsuitable for the scenarios we address [72] (more rationales in Appendix A-F).
@@
We selected the initial value $d$ as 267 days because, within this time frame, all four detectors remain optimal.
Detection results (Figure 13). As the time range was reduced, the performance of recall of the heuristic began to decline due to the limited number of order activities, which caused the heuristic validators to fail to trigger some detections. Given that the true negative count in the ground truth (316,049) is substantial compared to the true positive (3,117), both accuracy and precision of the heuristic remained unchanged when $d$ decreased. In contrast, the three ML algorithms maintained stable and good performances with decreasing time ranges. However, Random Forest performs consistently better than the other two ML algorithms in most $d$.
Random Forest maintained an accuracy greater than 93% down to $d$ = 56, whereas Logistic Regression and XGBoost saw their accuracy drop to 93% when using data within a time range of 58 days from the LP’s deployment date. Starting from $d$ = 56, the accuracy of all ML algorithms fell to 93%, marking the threshold at which all algorithms lost their high performance observed previously.
@@
VII Discussion
VII-A SLID Mitigation
@@
    Assess on-chain data. Prioritize transparency by thoroughly evaluating on-chain activities. Regularly monitor the tokens owned by LP owners for any signs of gradual drain.
VII-B Comparison to “Similar” Concepts
Conventional stock dilution. SLID scams closely resemble unauthorized stock issuance in traditional finance, where inflating the supply of shares without shareholder approval is a clear form of fraud. Regulatory bodies like U.S. Securities and Exchange Commission (SEC) enforce rules, which requires shareholder consent before issuing additional stock [75, 76]. A notable case is the SEC’s 2016 enforcement action against Goldman Sachs for facilitating naked short selling [75]. Similarly, in SLID, token issuers retain unilateral control over liquidity or token supply, diluting existing holdings without transparency or user consent - practices that would be considered illegal in traditional markets [77, 78].
Arbitrage bot. Those bots are automated agents designed to exploit price discrepancies across different markets or pools, often executing rapid trades to capture profit without taking directional market risk. Their behavior may involve frequent interactions with LPs, such as buying low in one and selling high in another, but they operate within transparent, rule-based systems and do not manipulate token supply or LP ownership. Unlike SLID scammers, arbitrage bots do not retain special privileges over token minting or LP tokens, and their profits stem from legitimate market inefficiencies rather than exploitative control. Therefore, despite their aggressive trading patterns, arbitrage bots are fundamentally different from SLID scams in both mechanism and intent.
Memecoins. We use the recently popular Trump token as a representative example of memecoins. While it shares superficial traits with SLID scams, such as rapid insider accumulation, it does not meet our criteria for SLID. In this case, insiders and affiliates acquired their tokens directly from the LP shortly after its launch (approximately two minutes post-deployment). Importantly, these tokens were not freely minted or allocated without cost. Instead, they were purchased at the prevailing market price, meaning that insiders assumed genuine financial risk and did not benefit from disproportionate or zero-cost supply advantages [79].
Their differences are evident from a financial perspective.
  • Asset origin: SLID scammers exploit the system by either minting tokens without disclosure or retaining their initial LP tokens, enabling them to later extract profits fraudulently at no cost. In contrast, legitimate traders acquire assets transparently at market prices and cannot mint unlimited tokens or withdraw the entire pool’s value.
  • Risk exposure: Legitimate traders bear genuine market risk, whereas SLID scammers face minimal risk due to the control over token minting and huge liquidity.
VII-C Adversarial Adaptation and Real-world Integration
Advarsarial adaptaion and trade-offs. While our detection methods are effective against canonical SLID patterns, adaptive adversaries may attempt to evade detection by modifying their extraction strategy. One plausible approach is to slow down profit-taking activities or reduce the number of withdrawal or sell actions below our heuristic thresholds (e.g., fewer than five profit-taking orders or extended delay beyond 60 days). Although this behavior may temporarily evade our rule-based validator, it introduces a critical trade-off: the scammer must significantly reduce their short-term return on investment (ROI) and prolong their attack window, increasing the risk of manual discovery or external scrutiny. In contrast, our machine learning model, trained on early-stage LP activity, remains effective even when profit-taking is delayed. This is because it captures structural anomalies such as retained LP token control, disproportionate minting of paired tokens, lack of LP token burning, and abnormal trading volume patterns. Thus, delaying extraction may reduce signal strength marginally but does not meaningfully bypass ML detection within the first 57 days.
Another adaptive strategy involves distributing fraudulent actions across multiple EOAs or using a different address from the deployer, thereby masking the scammer’s identity and mimicking legitimate user behavior. While this tactic may degrade the performance of purely wallet-based heuristics, the aggregate behavioral patterns (e.g., centralized token control, orchestrated sell timing, and highly correlated transaction profiles) remain detectable by our ML classifier. We emphasize that both heuristic and ML models exhibit adversarial robustness by design: the former requires multiple observable actions to trigger, while the latter generalizes across behavioral patterns rather than specific identities. Nonetheless, we recognize the value of formal adversarial resilience testing. Future work could quantify the trade-off between scam profitability and detection delay. This analysis would further highlight the dilemma faced by SLID scammers: maximizing stealth often requires sacrificing profitability or operational efficiency.
Real-world integration. Our solution can integrate directly with Ethereum DEXs as an LP monitoring and early warning service, providing real-time alerts on suspicious liquidity pools and enabling proactive fraud detection. By extracting key features from token transaction histories, it continuously tracks evolving on-chain behaviors while scaling efficiently to handle diverse market conditions. In addition, integration with payment service providers allows our system to evaluate wallet credibility by identifying addresses linked to SLID creation and analyzing suspicious transaction histories, even when legitimate tokens like Bitcoin are involved. This combined approach enhances risk assessment, blocks malicious token flows before they affect end users, and supports the development of stronger DeFi security standards.
VIII Conclusion